# Mobile Edge Intelligence for Large Language Models: A Contemporary Survey

### v1

@@ VII-C Split Learning
SL allows an edge server to take over the major training load from edge devices based on model splitting [214]. Since its invention, SL has been applied to various scenarios, such as healthcare [215, 216]. Unlike FL, SL only places a sub-model on edge devices for training, thereby considerably decreasing the workload of edge devices. The vanilla SL involves sequential interactions between an edge server and edge devices, which is a major bottleneck due to the waiting time of idle edge devices. The variants of SL, including parallel split learning (PSL) [217, 218] and SFL [219, 220], can be applied to enable parallel training while utilizing the resources on multiple edge devices. Clearly, analogous to FL, SL can also be integrated with PEFT or other resource-efficient techniques in Section III to alleviate the workload on edge devices further. For instance, edge devices are only required to execute the forward pass by freezing client-side parameters, considerably reducing the computing workload and memory usage.
@@ VIII Edge Inference for LLMs
Edge LLM inference leverages edge computing to provide inference services to end users based on well-trained LLMs. The major difference between edge inference for LLMs and conventional edge inference framework lies in the exploitation of LLM properties, such as multimodality, parameter-shareability, and the autoregressive process, to accelerate the inference process. Edge LLM inference can be categorized into three frameworks: centralized edge inference, split inference, and collaborative inference, as illustrated in Fig. 13. In this section, we will explore the associated challenges and review the efficient techniques to solve them. The related works on edge LLM inference are summarized in Table V.
(a) Centralized edge inference.
(b) Split inference.
(c) Collaborative inference.
Figure 13: Edge inference frameworks for LLMs. This figure presents the associated challenges and corresponding techniques in each sub-figure. In Fig. 13(a), edge devices centralized the data to the edge server with LLMs for centralized edge inference. In Fig. 13(b), edge devices feed the data into the user-side sub-LLM and upload the intermediate features to the edge server. Subsequently, the edge server conducts the inference with the server-side sub-LLM and returns the inference results to edge devices. In Fig. 13(c), edge devices first generate preliminary inference results with the user-side lightweight LLMs and upload the results to the edge server with the larger LLM for verification.
TABLE V: Summary of related works on edge LLM inference.
@@ VIII-A Centralized Edge Inference
In centralized edge LLM inference, edge devices offload their input token to edge servers for AI inference [16]. Note that efficient LLM inference methods, mentioned in Section III-A, can be implemented on edge servers to accelerate edge LLM inference. Centralized edge inference involves raw data uploading from edge devices to edge servers, which shares similar challenges with centralized edge learning in Section VII-A, i.e., communication latency for uploading data and computing burden (including excessive memory usage) on edge servers. In what follows, we will elaborate on the techniques to overcome these challenges.
@@ VIII-A3 Resource management for centralized LLM inference
Furthermore, one particular challenge in edge LLM inference lies in the autoregressive process of typical LLMs. This characteristic introduces two kinds of difficulty: 1) the task computing delay can hardly be predicted since the length of output tokens during the decoding process may not be known at the beginning, and 2) the memory usage is relevant to the length of output tokens according to KV cache, implying that we must reserve sufficient memory and computing resources for task scheduling. Based on the above observations, we can explore new mathematical modeling of edge computing to address resource optimization for edge LLM inference.
@@ VIII-B1 Split inference for LLMs with token representation reduction
Considering split inference, there are various ways to reduce the volume of token representations (latent representation) at the cut layer in LLMs for uploading to edge servers. First, the outputs of intermediate layers in Transformers can be compressed via quantization [249, 250, 115], pruning [234, 238], and merging [251, 252]: Uploading the compressed intermediate layer outputs can substantially reduce the communication overhead, thereby achieving low-latency split inference. 1) Quantization enables converting intermediate outputs of hidden layers in Transformers from high bitwidth to low bitwidth. In [233], the authors insert a binarization module after the layer norm layer inside the Transformer encoder blocks to represent the token representations with 1-bit vectors instead of float-value vectors. 2) Pruning involves removing redundant vectors of token representations, which can apply to both images and text inputs based on a lightweight module [238]. For example, the authors in [234] eliminate the redundant vectors of the intermediate encoder outputs, where the remaining intermediate encoder outputs of the last encoder only account for 2% of the original intermediate encoder outputs. With this method, edge devices can eliminate unimportant vectors of intermediate layer outputs before uploading them to edge servers in split LLM inference. As the size of remaining representation vectors decreases as the encoder goes deeper, which indicates that more encoder blocks remain on edge devices for local computing, it is worthwhile investigating the optimal cut layer selection involving the trade-off between communication efficiency and computing burden on edge devices. 3) Merging can compress the information of intermediate outputs by merging redundant token representations with similar semantic meanings. The authors in [235] propose ToMe inserted into the Transformer blocks, which merges the representations of similar tokens in the outputs of MHAs. It is promising to employ this method in split LLM inference so that edge devices merge the representations in intermediate results with the lightweight merging module before uploading to edge servers.
@@ VIII-B3 Split inference with early exit
Early exit techniques, as introduced in Section III-A, have been widely used in LLM inference for latency reduction. This approach can be adopted in split inference to reduce computing latency on edge devices/servers. In this context, when adding early exit modules into user-side sub-LLMs, the inference computation in later layers in the user(server)-side sub-LLMs can be skipped, implying that there is no need for uploading intermediate features for latency reduction. However, in LLM inference with the early exit approach, the hidden states of the corresponding token in the later layers are missing when the token outputs the results in the early exit layer. Therefore, in split LLM inference with the early exit technique, if the early exit layers reside at user-side sub-LLMs, the hidden states still need to be uploaded to edge servers for KV cache computation in later layers. As such, we need to design the uploading strategy of hidden states at exiting layers, e.g., opportunistically uploading the hidden states by considering channel conditions.
@@ VIII-C Collaborative inference
Split inference exchanges high-dimensional intermediate features, resulting in excessive communication overhead. To overcome this limitation, edge devices and edge servers can cooperate in some other modes with the smaller size of information exchange. For instance, speculative decoding [95], which has been introduced in Section III-A, can be adopted in device-server collaborative inference for LLMs. Speculative decoding enables an edge device to run a smaller on-device LLM, called an approximation model, while asking the edge server to run a larger LLM to verify and correct the output tokens uploaded by the edge device. The main advantages of this approach are threefold: First, it empowers edge devices to generate preliminary results/decisions, which can be used for low-latency inference. Second, the parallel decoding process accelerates the inference process on the edge server. Since LLMs generate tokens with the autoregressive decoding technique, verifying a long sequence of tokens is much faster than decoding serially by the server-side LLM alone. Third, concerning communication overhead, the output token can be smaller than cut-layer intermediate features in many cases. To further save communication-computing resources, edge devices can decide whether to upload the tokens generated by on-device LLMs to edge servers for verification [242] based on calibrated confidence scores, because highly confident outputs might not require an edge server to consume resources for verification.
@@ VIII-D Lessons Learned
Edge devices and servers usually cooperate to support LLM inference for privacy and latency concerns. Split or collaborative inference is constrained by various factors, including memory and communication-computing latency. Considering these impact factors, the optimization of edge LLM inference must be carefully investigated by considering various LLM techniques, such as KV cache optimization, multimodal feature extraction, and the autoregressive model. All of these characteristics significantly impact the memory usage, communication overhead, and inference latency for LLMs. For this reason, the design of edge LLM inference, such as model splitting, early exit, and resource allocation, naturally differs from conventional edge inference systems, creating a rich set of optimization challenges and opportunities in this area.
@@ IX Further Research Opportunities
MEI for LLMs remains a largely uncharted direction. This section will provide in-depth discussions on several pressing concerns and how we can address them.
@@ IX-A Green Edge LLM
Energy consumption is a primary public concern over LLMs. It is estimated that the energy consumption of training GPT-4 is equivalent to the consumption over 5 to 6 years of 1,000 average US households. Moreover, the energy costs of model inference can be even higher because of the frequent service requests from users worldwide. Therefore, one pressing research issue is designing energy-efficient LLM training and inference.
@@
Edge LLM can reduce the energy consumption of LLMs in three aspects. First, making model fine-tuning or inference available at the network edge eliminates the need to transmit large volumes of data to cloud centers, thereby reducing the energy costs in backbone networks. Second, integrated communication-computing design can be jointly optimized to enhance energy efficiency for training/inference. For instance, less data communication can be carried out by data compression or parameter freezing to reduce total transmit power as long as the required training/inference accuracy can be achieved. At last, edge LLM can utilize a small-scale LLM to obtain initial inference results and harness the power of the cloud-based large-scale LLM only when the inference confidence is low. This potentially reduces the energy consumption at cloud centers resulting from invoking large-scale LLMs for every user request.
@@
Regarding the research problems, green edge LLM features the integrated design of wireless communications and computing, which must account for the overall transmission and computing energy. With this in mind, there are two main design goals, i.e., reducing energy consumption on edge devices and decreasing the overall energy consumption for green AI. The first one benefits battery-constrained IoT and mobile devices, making AI services or training more accessible to customers. To achieve this goal, AI training/inference can be offloaded to edge servers as long as it saves energy. For instance, when adopting PEFT and SL, the client-side models can be frozen as much as possible, even entirely frozen, to minimize energy consumption on edge devices. Although this approach might lead to slower learning convergence or increase the energy consumption on edge servers for achieving a target training/inference accuracy, this cost might be less concerning since edge servers are usually more powerful and connected to powerlines. The second one aims to decrease the overall system costs, particularly from the perspective of mobile operators. By minimizing the total energy consumption or maximizing the overall energy efficiency, network operators can improve AI-centric metrics with limited or lower energy costs. A meaningful metric can be “AI energy efficiency” [257], i.e., the amount of intelligence (AI accuracy) achieved per energy cost under the constraints of basic accuracy requirements. The optimization of AI by taking energy consumption into consideration can eliminate the case where the system uses a great amount of energy to merely achieve minor improvements, which is not justifiable from both operators’ and society’s viewpoints.
@@ IX-B Secure Edge LLM
Secure edge LLM is another important branch of research. While LLM security has been extensively studied, edge LLM security has received much less attention. Specifically, edge LLM usually involves distributed learning of LLMs, where new challenges arise. Although federated learning and split learning, as alluded to earlier, are privacy-enhancing approaches for LLM to avoid direct access to personal data, there is still privacy risk as malicious servers may launch attacks to recover the raw data based on received models or intermediate features [258, 259]. At the mobile edge, tailored security components should be integrated to ensure secure and robust LLMs.
@@
Let us consider two aspects of secure edge LLM. The first aspect is defending against inference attacks to protect user privacy leakage from mobile edge servers or other edge devices. It has been demonstrated that prompting LLMs may reveal the private information of other users, such as credit card information, in the training process by only inserting a few benign-appearing sentences into the training dataset [260]. A similar problem may also be observed in FL and SL, creating a rich set of research problems. One potential defense mechanism is to add noise to highly sensitive personal data, such as credit card information, based on the theory of differential privacy. Also, appropriate mechanisms may be developed to detect such benign-appearing tokens/parameters/features from clients.
@@
The second aspect is defending against data poisoning or backdoor attacks by filtering out malicious users who aim to alter the training process, thereby maintaining training effectiveness. These attacks can lead to severe consequences for LLMs, i.e., outputting harmful health guidance if considering a healthcare LLM. Although such attacks have been studied for LLMs [261, 262], there is still a lack of research works on designing effective attack/defense mechanisms by considering distributed learning for LLMs, particularly FL and SL, over edge devices. Considering PEFT, the attacker may only be able to change a small proportion of the models, e.g., adaptors or prompts, to impact the training process, leading to new challenges/opportunities in designing the attack/defense schemes.

### v2

@@ VII-C Split Learning
SL allows an edge server to take over the major training load from edge devices based on model splitting [250]. Since its invention, SL has been applied to various scenarios, such as healthcare [251, 252]. Unlike FL, SL only places a sub-model on edge devices for training, thereby considerably decreasing the workload of edge devices. The vanilla SL involves sequential interactions between an edge server and edge devices, which is a major bottleneck due to the waiting time of idle edge devices. The variants of SL, including parallel split learning (PSL) [253, 254] and SFL [255, 256], can be applied to enable parallel training while utilizing the resources on multiple edge devices. Clearly, analogous to FL, SL can also be integrated with PEFT or other resource-efficient techniques in Section IV to alleviate the workload on edge devices further. For instance, edge devices are only required to execute the forward pass by freezing client-side parameters, considerably reducing the computing workload and memory usage.
@@ VIII Edge Inference for LLMs
Edge LLM inference leverages edge computing to provide inference services to end users based on well-trained LLMs. Inference services in edge networks involve processing input data, such as text, images, and audio, to generate results like classifications, predictions, and responses with AI models for edge devices, including mobile users and IoT devices. One example is LLM-empowered robotic applications, where a robot uploads captured audio, video, and radar/lidar data across different modalities for feeding into a well-trained LLM on an edge server, thus facilitating the robot to understand the environments and execute actionable policies. The major difference between edge inference for LLMs and conventional edge inference framework lies in the exploitation of LLM properties, such as multimodality, parameter-shareability, and the autoregressive process, to accelerate the inference process. As illustrated in Fig. LABEL:fig:edge_inference, this section elaborates edge LLM inference through three categories: centralized edge inference, split inference, and collaborative inference, with the related works summarized in Table VI.
TABLE VI: Summary of related works on edge LLM inference.
@@ VIII-A Centralized Edge Inference
In centralized edge LLM inference, edge devices offload their input token to edge servers for AI inference [24]. Note that efficient LLM inference methods, mentioned in Section IV-A, can be implemented on edge servers to accelerate edge LLM inference. Centralized edge inference involves raw data uploading from edge devices to edge servers, which shares similar challenges with centralized edge learning in Section VII-A, i.e., communication latency for uploading data and computing burden (including excessive memory usage) on edge servers. In what follows, we will elaborate on the techniques to overcome these challenges.
@@ VIII-A3 Resource management for centralized LLM inference
Furthermore, one particular challenge in edge LLM inference lies in the autoregressive process of LLMs. This characteristic introduces two kinds of difficulty: 1) the task computing delay can hardly be predicted since the length of output tokens during the decoding process may not be known at the beginning, and 2) the memory usage is relevant to the length of output tokens according to KV cache, implying that we must reserve sufficient memory and computing resources for task scheduling. Based on the above observations, we can explore new mathematical modeling of edge computing to address resource optimization for edge LLM inference.
@@ VIII-B1 Split inference for LLMs with token representation reduction
Considering split inference, there are various ways to reduce the volume of token representations (latent representation) at the cut layer in LLMs for uploading to edge servers. First, the outputs of intermediate layers in Transformers can be compressed via quantization [284, 285, 156], pruning [270, 274], and merging [286, 287]. Uploading the compressed intermediate layer outputs can substantially reduce the communication overhead, thereby achieving low-latency split inference. 1) Quantization enables converting intermediate outputs of hidden layers in Transformers from high bitwidth to low bitwidth. In [269], the authors insert a binarization module after the layer norm layer inside the Transformer encoder blocks to represent the token representations with 1-bit vectors instead of float-value vectors. 2) Pruning involves removing redundant vectors of token representations, which can apply to both images and text inputs based on a lightweight module [274]. For example, the authors in [270] eliminate the redundant vectors of the intermediate encoder outputs, where the remaining intermediate encoder outputs of the last encoder only account for 2% of the original intermediate encoder outputs. With this method, edge devices can eliminate unimportant vectors of intermediate layer outputs before uploading them to edge servers in split LLM inference. As the size of remaining representation vectors decreases as the encoder goes deeper, which indicates that more encoder blocks remain on edge devices for local computing, it is worthwhile investigating the optimal cut layer selection involving the trade-off between communication efficiency and computing burden on edge devices. 3) Merging can compress the information of intermediate outputs by merging redundant token representations with similar semantic meanings. The authors in [271] propose ToMe inserted into the Transformer blocks, which merges the representations of similar tokens in the outputs of MHAs. It is promising to employ this method in split LLM inference so that edge devices merge the representations in intermediate results with the lightweight merging module before uploading to edge servers.
@@ VIII-B3 Split inference with early exit
Early exit techniques, as introduced in Section IV-A, have been widely used in LLM inference for latency reduction. This approach can be adopted in split inference to reduce computing latency on edge devices/servers. In this context, when adding early exit modules into user-side sub-LLMs, the inference computation in later layers in the user(server)-side sub-LLMs can be skipped, implying that there is no need for uploading intermediate features for latency reduction. However, in LLM inference with the early exit approach, the hidden states of the corresponding token in the later layers are missing when the token outputs the results in the early exit layer. Therefore, in split LLM inference with the early exit technique, if the early exit layers reside at user-side sub-LLMs, the hidden states still need to be uploaded to edge servers for KV cache computation in later layers. As such, we need to design the uploading strategy of hidden states at exiting layers, e.g., opportunistically uploading the hidden states by considering channel conditions.
@@ VIII-C Collaborative Inference
Split inference exchanges high-dimensional intermediate features, resulting in excessive communication overhead. To overcome this limitation, edge devices and edge servers can cooperate in some other modes with the smaller size of information exchange. For instance, speculative decoding [136], which has been introduced in Section IV-A, can be adopted in device-server collaborative inference for LLMs. Speculative decoding enables an edge device to run a smaller on-device LLM, called an approximation model, while asking the edge server to run a larger LLM to verify and correct the output tokens uploaded by the edge device. The main advantages of this approach are threefold: First, it empowers edge devices to generate preliminary results/decisions, which can be used for low-latency inference. Second, the parallel decoding process accelerates the inference process on the edge server. Since LLMs generate tokens with the autoregressive decoding technique, verifying a long sequence of tokens is much faster than decoding serially by the server-side LLM alone. Third, concerning communication overhead, the output token can be smaller than cut-layer intermediate features in many cases. To further save communication-computing resources, edge devices can decide whether to upload the tokens generated by on-device LLMs to edge servers for verification [278] based on calibrated confidence scores, because highly confident outputs might not require an edge server to consume resources for verification.
@@ VIII-D Lessons Learned
Resource-efficient LLM inference techniques can be adapted to wireless edge environments, enabling effective cooperation between edge devices and edge servers. Thus, edge LLM inference must be jointly optimized with various LLM techniques, such as KV cache optimization, multimodal feature extraction, and autoregressive decoding, all of which significantly impact memory usage, communication overhead, and computing latency for LLM inference. For this reason, the design of edge LLM inference, such as model splitting, early exiting, and radio-computing resource allocation, naturally differs from conventional edge inference systems, creating a rich set of research problems in this area.
@@ IX Further Research Opportunities
MEI for LLMs remains a largely uncharted direction. This section will discuss the pressing concerns and potential solutions for MEI4LLM.
@@ IX-A Green Edge LLM
Energy consumption is a primary public concern over LLMs. For instance, it is estimated that the energy consumption of training GPT-4 is equivalent to the consumption over 5 to 6 years of 1,000 average US households. Moreover, the energy costs of model inference can be even higher due to the frequent service requests from users worldwide. The estimated energy footprint of GPT-4 training and inference are detailed in Table VII. These facts underscore the urgent need for the development of energy-efficient LLM training and inference techniques.
TABLE VII: Approximate energy usage of GPT models.
Name Total energy consumption in training (MWh) Daily energy consumption of responses (MWh)
GPT-3 1,287 [292] 3 [293]
GPT-4 $\geq$50,000 [294] 5 [293]
@@
By pushing the intelligence to the network edge, MEI4LLM can reduce the energy consumption of LLMs in three aspects. First, making model fine-tuning or inference services available at the network edge eliminates the need to transmit large volumes of data to cloud centers, thereby reducing the energy costs in data transportation. Second, integrated communication-computing design can be jointly optimized to enhance energy efficiency for training/inference. For instance, less data communication can be achieved by data compression or parameter freezing to reduce total transmit power as long as the required training/inference accuracy can be achieved. At last, edge LLM can utilize a small-scale LLM to obtain initial inference results and harness the power of the cloud-based large-scale LLM only when the inference confidence is low. This potentially reduces the energy consumption at cloud centers resulting from invoking large-scale LLMs for every user request.
@@
Regarding the research problems, green edge LLM features the integrated design of wireless communications and computing, which must account for the overall transmission and computing energy. With this in mind, there are two main design goals, i.e., reducing energy consumption on edge devices or decreasing the overall energy consumption for green AI. The first goal is to benefit battery-constrained IoT and mobile devices, making AI services or training more accessible to customers. To achieve this goal, AI training/inference can be offloaded to edge servers as long as it saves energy for end users. For instance, when adopting PEFT and SL, the client-side models can be frozen as much as possible, even entirely frozen, to minimize energy consumption on edge devices. Although this approach might lead to slower learning convergence or increase the energy consumption on edge servers for achieving a target training/inference accuracy, this cost might be less concerning since edge servers are usually more powerful and connected to powerlines. The second one aims to decrease the overall system costs, particularly from the perspective of mobile operators. By minimizing the total energy consumption or maximizing the overall energy efficiency, network operators can improve AI-centric metrics with limited or lower energy costs. A meaningful metric can be “AI energy efficiency” [295], i.e., the amount of intelligence (AI accuracy) achieved per energy cost. The optimization of AI by considering energy consumption can eliminate the case where the system uses a great amount of energy to merely achieve minor improvements, which is not justifiable from both operators’ and society’s viewpoints.
@@ IX-B Secure Edge LLM
Secure edge LLM is another important branch of research. Although federated learning and split learning, as alluded to earlier, can serve as privacy-enhancing training frameworks for LLMs to avoid direct access to personal data, there is still privacy risk as malicious servers may launch attacks to recover the raw data based on received models or intermediate features [296, 297]. However, while LLM security has been extensively studied, edge LLM security has received much less attention. At the mobile edge, security components should be integrated to ensure secure and trustworthy LLMs.
@@
Let us consider two aspects of secure edge LLM. The first aspect is defending against inference attacks to protect user privacy leakage from honest-but-curious edge servers or other edge devices. It has been demonstrated that prompting LLMs may reveal the private information of other users, such as credit card information, in the training process by only inserting a few benign-appearing sentences into the training dataset [298]. For FL and SL, similar problems may still exist. From the user perspective, one potential defense mechanism is to add noise to protect highly sensitive personal data, such as credit card information, based on the theory of differential privacy. The research question is where to insert noise and the amount of noise to insert.
@@
The second aspect is defending against data poisoning or backdoor attacks by filtering out malicious users who aim to alter the training process, thereby maintaining training effectiveness. These attacks can lead to severe consequences for LLMs, i.e., outputting harmful health guidance if considering a healthcare LLM. Although such attacks have been studied for LLMs [299, 300], there is still a lack of research works on designing effective attack/defense mechanisms by considering distributed learning for LLMs, particularly FL and SL, over edge devices. Considering PEFT, the attacker may only be able to change a small proportion of the models, e.g., adaptors or prompts, to impact the training process rather than modifying the full set of parameters. This leads to new challenges/opportunities in designing the attack/defense schemes.
IX-C Quality-aware Edge LLM Training
Unlike cloud-based training, data quality at the network edge cannot be well controlled due to the lack of human annotations and monitoring. Unfortunately, training over massive and low-quality data/labels can be detrimental to model performance rather than beneficial. To effectively deploy LLMs at the network edge, it is essential to enable LLM training with automated quality control.
There are two key research directions for improving the quality of edge LLM training. First, improving the quality of training datasets can directly improve edge LLM training outcomes. On the one hand, edge servers can increase the diversity of training datasets by prioritizing varied data sources. Higher diversity in training datasets increases the likelihood of achieving better performance on corresponding tasks [301]. For example, in federated learning, edge servers can allocate additional training rounds to edge devices with more diverse data. Additionally, in cloud-edge-end FL, the cloud center can assign varying weights to edge LLMs during aggregation based on the regional diversity associated with different edge servers. On the other hand, to ensure high accuracy and reliability in LLM responses, data cleaning for training datasets can be adopted for edge LLM training. Considering LLMs’ sensitivity to input data, data cleaning is essential to detect and remove errors, irrelevant information, duplicates, and inconsistencies before LLM training [302]. For example, in LLM-empowered AI chatbots, data cleaning can identify and remove useless information and grammatical errors in raw data inputs, thereby enhancing the quality of the training datasets and ensuring LLMs’ performance [303]. However, since data cleaning introduces additional computational overhead to edge servers/devices, a trade-off exists between computational efficiency and the quality of training data.
Second, edge LLM training can harness a large volume of unlabeled data or data with noisy labels captured by edge devices. While the training quality of LLMs significantly depends on high-quality training datasets, acquiring enough high-quality labeled data for edge devices requires considerable effort and costs, particularly for the vast number of resource-limited IoT devices at the network edge. Therefore, edge servers must leverage the large volume of unlabeled data collected from distributed edge networks to train LLMs with unsupervised learning [304] and semi-supervised learning. For LLMs, edge servers can fully leverage the relationships between limited labeled data and large volumes of unlabeled data across different modalities. For example, on-device LLM can continuously capture data from multiple modalities, automatically generating labels for data of specific modalities. By utilizing the inherent patterns and correlations in these data, edge servers can reconstruct the labels of unlabeled data, thereby enhancing the training quality and adaptability across various tasks.